OpenStudy (anonymous):

find the radian measure that corresponds to the given degree measure. b) 750 degree c) -135 degree d) -330 degree

OpenStudy (anonymous):

Hint: 180 degree = pi radian

OpenStudy (anonymous):

what do 750 and 180 have in common? @sauravshakya

OpenStudy (anonymous):

180 degree = pi radian 1 degree = pi/180 radian 750 degree = 750 * pi /180 radian =13.09 radian

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

Can u do the rest